英語 (English)

基盤英語 (Basic English) (総科 (全 1 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 1 単位前期木 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】This class will focus on acquiring the proficiency in English to make further study in any field at college. Acquiring the knowledge of a foreign language and the proficiency in it is to This classs designed to give a good assistance in developing Basic English abilities in general. Especially, It will focus on Grammar, Listening, Reading and Writing skills.

【授業の概要】This class is organized to help you improve your basic English ability covering skills like reading, writing and listening, which will surely help you try TOEIC and other tests.

【キーワード】 TOEIC， Structure of English Language， General English， Writing 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. To acquire the knowledge of the structure of English language through writing.

2. To find an appropriate key to improve English skills. 【授業の計画】

1. Overall Guidance

2. Clip 1

3. Clip 2

4. Clip 3

5. Clip 4

6. Clip 5

7. Clip 6

8. Clip 7

9. Clip 8

10. Clip 9

11. Clip 10

12. Clip 11

13. Clip 12

14. Clip 13

15. Overview 【教科書】 No specific text book but there will be some other media will be selected

up to date. 【参考書等】 References and other aid materials will be introduced and prepared in

class. 【成績評価の方法】 Evaluation is made considering your fuul attendance,good

participation and homework sheet only. 【再試験の有無】 No make-up examination for this course. 【受講者へのメッセージ】日々の積み重ねを重視してください．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220821 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 座喜 . (オフィスアワー: contact E-mail: jzaki@ias.tokushima-u.ac.jp)

English

Basic English (総科 (全 1 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 1 unit 前期木 1 ・ 2

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ This class will focus on acquiring the proficiency in English to make further study in any field at college. Acquiring the knowledge of a foreign language and the proficiency in it is to This classs designed to give a good assistance in developing Basic English abilities in general. Especially, It will focus on Grammar, Listening, Reading and Writing skills.

Outline⟩ This class is organized to help you improve your basic English ability covering skills like reading, writing and listening, which will surely help you try TOEIC and other tests.

�������⟩ TOEIC, Structure of English Language, General English, Writing Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. To acquire the knowledge of the structure of English language through writing.

2. To find an appropriate key to improve English skills.

Schedule⟩

1. Overall Guidance

2. Clip 1

3. Clip 2

4. Clip 3

5. Clip 4

6. Clip 5

7. Clip 6

8. Clip 7

9. Clip 8

10. Clip 9

11. Clip 10

12. Clip 11

13. Clip 12

14. Clip 13

15. Overview Textbook⟩ No specific text book but there will be some other media will be

selected up to date. Reference⟩ References and other aid materials will be introduced and prepared in

class. Evaluation Criteria⟩ Evaluation is made considering your fuul attendance,good

participation and homework sheet only. Re-evaluation⟩ No make-up examination for this course. Message⟩ Practice makes perfect: study everyday.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220821 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki . (Office Hour: contact E-mail: jzaki@ias.tokushima-u.ac.jp)